OpenStudy (kaleidoscopicsink):

Find the slope of the graph of the relation x^2y + 4y = 8 at the point (2, 1). I keep getting -1/8 and I'm really confused. Any help is accepted and greatly cherished.

OpenStudy (tkhunny):

How do you manage that? Please demonstrate your process.

OpenStudy (kaleidoscopicsink):

4y+x^2y --> 4 (dy/dy) + (y+ 2x * dy/dx) --> 4 (dy/dy) +2x(dy/dx) = -y --> -y / 4+2x = dy/dx **Which I know is horribly wrong

OpenStudy (tkhunny):

?? "Which I know is horribly wrong" -- Then why did you do it? [x^2y] + 4y = 8 Let's take a closer look at the multiplication rule. [x^2 (dy/dx) + 2xy] + 4 (dy/dx) = 0

OpenStudy (kaleidoscopicsink):

Okay, give me a second, I forgot about the product rule.

OpenStudy (kaleidoscopicsink):

Would the answer be -1/2 for the slope?

OpenStudy (tkhunny):

Again, how did you get that?

OpenStudy (tkhunny):

[x^2 (dy/dx) + 2xy] + 4 (dy/dx) = 0 (2,1) [2^2 (dy/dx) + 2(2)(1)] + 4 (dy/dx) = 0 4 (dy/dx) + 4 + 4 (dy/dx) = 0 dy/dx = -4/8 = -1/2 Show your work in the future, please. It is so much easier to redirect your from errors and encourage you with acceptable results.

OpenStudy (kaleidoscopicsink):

Using the product rule for x^2y you get: 2xy+x^2. Then going back to 4y being 4 --> 4 (dy/dx) + 2xy + x^2 * dy/dx --> -2xy / 4+x^2 then insert 1,2.

OpenStudy (tkhunny):

Pretty good, except for the tiny notation problem. You mean -2xy / [4 + x^2]. You do NOT mean what you have written.

OpenStudy (kaleidoscopicsink):

Yep. cx Can you check if I did this right? If 5x^2 + y^4 = −9 then evaluate the second derivative of y with respect to x when x = 2 and y = 1 So 5x^2 goes to 10x and y^4 goes to 4y^3 then that derivative is -10x/4y^3, then take the derivative of that using the quotient rule which gives you -40y^3 +120xy^2 / 16y^6, finally insert the given vaalues and you get 12.5? @tkhunny

OpenStudy (tkhunny):

Please post on another thread. Sorry, gtg.

OpenStudy (kaleidoscopicsink):

It's okay. Have a great day. Thanks for the help. cx @tkhunny